# Towards Reliable Proof Generation with LLMs: A Neuro-Symbolic Approach

### v3

@@ 1 Introduction
Figure 2: An example problem from FormalGeo-7k (level 3). Left: Problem inputs, including a natural language description and formal representations – construction (entities and relations), extended construction (inferred based on extension rules), conditions, and goal. The input also includes a numeric answer (or expression) and a formal proof, composed of steps, each invoking a theorem from the dictionary. Right: The Theorem Dictionary. Theorems include variation id, arguments, premises and conclusions. The full dataset also includes diagrams, but we do not use them here. A recent work Zhang et al. (2024) has shown that multimodal LLMs often struggle with the visual aspects of math problems, often performing better without the images.
@@
Mathematical proofs demand rigorous logical deduction, symbolic manipulation, and an understanding of abstract concepts that go beyond statistical correlations in language. The requirement for absolute truth and the absence of ambiguity in mathematical reasoning present a significant challenge for models trained to generate plausible text rather than formally valid inferences Singh et al. (2024); Pan et al. (2025). In addition, the often lengthy nature of proofs necessitates a level of sustained logical coherence and hierarchical reasoning that is hard for current LLMs.
@@
As a proof of concept, we focus on Euclidean geometry – a domain that is symbolic, verifiable, and rich in structural analogies. Note that our primary objective in this paper is to evaluate whether our symbolic augmentations can improve the proof-generation capabilities of general-purpose LLMs. As such, while geometry is the domain we test our ideas in, our focus is not on competing with state-of-the-art, specialized geometry solvers. Rather, we wish to quantify the gains enabled by our method.
Our main contributions are:
@@
  • We design a symbolic verifier tailored to geometry proofs with expressive feedback.
@@
  • Our method reduces costs through focused context construction. Reducing the theorem dictionary from 17,950 tokens to just 2,513 tokens on average.
  • Our code and data are publicly available at our project website: https://www.orensultan.com/llm_proof_generator.
@@ 3.2 Analogous Problems Retrieval
The dataset contains $\sim 24.4$M problem pairs (all combinations of 6,981 problems). Proof similarity scores are binned into five intervals of width 0.2. As the distribution is highly imbalanced, with 88% of pairs in the first bin (0-0.2), we down-sample all bins to the size of the smallest (131K, 0.8-1), yielding a balanced dataset of 655K pairs. We split the data into 90% training and 10% evaluation.
@@ 3.3 LLM Proof Generation
Table 1: Average number of problems whose entire proof was covered by the union of theorems from their top-$k$ analogies vs. random problems, over 100 random target problems (20 per level, 1–5). Analogies consistently achieve higher coverage despite fewer theorems.
@@
To evaluate the effectiveness of this approach, we measure the extent to which the narrowed-down dictionary still captures all the theorems needed for the proof of the target problem. We sample 100 problems (20 from each level 1-5) and evaluate how many of their target proofs are completely covered using theorems from their top-$k$ similar problems (retrieved with the regressor of Section 3.2), and compare to a random set of $k$ problems.
@@ 6 Related Work
In another work, Zhou et al. (2025) approach analogical reasoning as a self-supervised learning task, fine-tuning models to transfer solutions between structurally similar problems. The model is trained to solve a target problem conditioned on the solution to a related source problem, learning abstract transformation patterns from many pairs. In contrast, our method requires no additional training and operates entirely at inference time.

### v4

@@ 1 Introduction
Figure 2: An example problem from FormalGeo-7k. Left: Problem inputs, including a natural language description and formal representations – construction (entities and relations), extended construction (inferred based on extension rules), conditions, and goal. The input also includes a numeric answer (or expression) and a formal proof, composed of steps, each invoking a theorem from the dictionary. Right: The Theorem Dictionary. Theorems include variation id, arguments, premises and conclusions. The full dataset also includes diagrams, but we do not use them, as a recent work Zhang et al. (2024) showed that multimodal LLMs often struggle with the visual aspects of math problems.
@@
Mathematical proofs demand rigorous logical deduction, symbolic manipulation, and an understanding of abstract concepts that go beyond statistical correlations in language. The requirement for absolute truth and the absence of ambiguity in mathematical reasoning present a significant challenge for models trained to generate plausible text rather than formally valid inferences Singh et al. (2024); Pan et al. (2025).
In addition, the often lengthy nature of proofs necessitates a level of sustained logical coherence and hierarchical reasoning that is hard for current LLMs. Recent work has shown that altering even superficial aspects of mathematical problems results in significant performance drops Mirzadeh et al. (2024), suggesting that their success often hinges on pattern matching rather than genuine mathematical reasoning.
@@
As a proof of concept, we focus on Euclidean geometry – a domain that is symbolic, verifiable, and rich in structural analogies. Note that our primary objective in this paper is to evaluate whether our symbolic augmentations can improve the proof-generation capabilities of general-purpose LLMs. As such, while geometry is the domain we test our ideas in, our focus is not on competing with state-of-the-art, specialized geometry solvers. Rather, we wish to quantify the gains enabled by our method. Our main contributions are:
@@
  • We design a symbolic verifier tailored to geometry proofs with expressive feedback. In contrast to other works, we evaluate the entire proof, not just the final numeric answer.
@@
  • Our method reduces costs via focused context construction, reducing the theorem dictionary from 18K to just 2.5K tokens on average.
  • We will release code and data, including the evaluation scripts and processed data used in our experiments [^1] URL redacted, will be available upon publication..
@@ 3.2 Analogous Problems Retrieval
The dataset contains 24.4$\sim$ 24.4∼ 24.4M problem pairs (all combinations of 6,981 problems). Proof similarity scores are binned into five intervals of width 0.2. As the distribution is highly imbalanced, with 88% of pairs in the first bin (0-0.2), we down-sample all bins to the size of the smallest (131K, 0.8-1), yielding a balanced dataset of 655K pairs. We split the data into 90% training and 10% evaluation.
@@ 3.3 LLM Proof Generation
[…]
@@
[…]
@@ 6 Related Work
Zhou et al. (2025) approach analogical reasoning as a self-supervised learning task, fine-tuning models to transfer solutions between structurally similar problems. The model is trained to solve a target problem conditioned on the solution to a related source problem, learning abstract transformation patterns from many pairs. In contrast, our method requires no additional training and operates entirely at inference time.
In another work, Lee et al. (2025) generate in-context examples for math word problems by creating numerical variants of the target problem. These examples are automatically created, unlabeled, and include only the problems, without solutions. In contrast, our focus is on generating formal mathematical proofs. Thus, in our setting altering numerical values offers virtually no benefits.